50 ml of 0.3M HNO3 is added to 50ml of 0.3M NH3 calculate ph
Answer #1

millimoels of HNO3 = 50 x 0.3 = 15

millimoles of NH3 = 50 x 0.3 = 15

here millimoles of acid and base are same. so salt remians here.

salt concentration = 15 / (50 + 50) = 0.15 M

pH = 7 - 1/2 (pKb + log C)

    = 7 - 1/2 (4.74 + log 0.15)

   = 5.04

pH = 5.04